A solid piece of lead has a mass of 36.98 g and a volume of 3.30 cm3. From these data, calculate the density of lead in SI units (kilograms per cubic meter)
1
Expert's answer
2020-05-18T10:27:07-0400

We can find the density of lead in SI units from the formula:


"\rho = \dfrac{m}{V} = \dfrac{36.98g \cdot \dfrac{1kg}{1000g}}{3.30 cm^3 \cdot \dfrac{1m^3}{10^6 cm^3}} = 11206 \dfrac{kg}{m^3}."

Answer:

"\rho = 11206 \dfrac{kg}{m^3}."